OpenStudy (anonymous):

solve the equation: e^(squr)(t)=x^4 for 't'

OpenStudy (anonymous):

introduce ln on both sides : ln[e^(sqrt(t))] = ln(x^4) sqrt(t) * ln[e] = 4ln(x) sqrt(t) = 4ln(x) t = 16 ln^2(x)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw